The controller of Bridgeport Housewares Inc. instructs you to prepare a monthly cash budget for the next three months. You are presented with the following budget information:

September October November
Sales $142,000 $169,000 $224,000
Manufacturing costs 60,000 73,000 81,000
Selling and administrative expenses 50,000 51,000 85,000
Capital expenditures _ _ 54,000

The company expects to sell about 10% of its merchandise for cash. Of sales on account, 70% are expected to be collected in the month following the sale and the remainder the following month (second month following sale). Depreciation, insurance, and property tax expense represent $7,000 of the estimated monthly manufacturing costs. The annual insurance premium is paid in January, and the annual property taxes are paid in December. Of the remainder of the manufacturing costs, 80% are expected to be paid in the month in which they are incurred and the balance in the following month.

Current assets as of September 1 include cash of $54,000, marketable securities of $77,000, and accounts receivable of $158,200 ($124,000 from July sales and $34,200 from August sales). Sales on account for July and August were $114,000 and $124,000, respectively. Current liabilities as of September 1 include $7,000 of accounts payable incurred in August for manufacturing costs. All selling and administrative expenses are paid in cash in the period they are incurred. An estimated income tax payment of $20,000 will be made in October. Bridgeport’s regular quarterly dividend of $7,000 is expected to be declared in October and paid in November. Management desires to maintain a minimum cash balance of $53,000.

Required:

1. Prepare a monthly cash budget and supporting schedules for September, October, and November. Input all amounts as positive values except overall cash decrease and deficiency which should be indicated with a minus sign. Assume 360 days per year for interest calculations.
Answer #1

Bridgeport Housewares Inc. Sales Budget Cash Credit Total Sales September October November 14200 16900 22400 127800 152100 20

1 Bridgeport Housewares Inc. 3 Sales Budget 5 Cash 6 Credit 7 Total Sales September =B7*0.1 =B7-B5 142000 October =C7*0.1 C7-